### v2

# Leveraging differentiable programming in the inverse problem of neutron stars

Abstract
[…]
NS probe the high-density regime of the nuclear equation of state (EOS). However, inferring the EOS from observations of neutron stars is a computationally challenging task. In this work, we efficiently solve this inverse problem by leveraging differential programming in two ways. First, we enable full Bayesian inference in under one hour of wall time on a GPU by using gradient-based samplers, without requiring pre-trained machine learning emulators. Moreover, we demonstrate efficient scaling to high-dimensional parameter spaces. Second, we introduce a novel gradient-based optimization scheme that recovers the EOS of a given NS mass-radius curve. We demonstrate how our framework can reveal consistencies or tensions between nuclear physics and astrophysics. First, we show how the breakdown density of a metamodel description of the EOS can be determined from NS observations. Second, we demonstrate how degeneracies in EOS modeling using nuclear empirical parameters can influence the inverse problem during gradient-based optimization. Looking ahead, our approach opens up new theoretical studies of the relation between NS properties and the EOS, while effectively tackling the data analysis challenges brought by future detectors.
preprint: LA-UR-25-23486
I Introduction
The nuclear EOS remains uncertain and an open problem in physics Lattimer:2012nd; Ozel:2016oaf; Burgio:2021vgk; Chatziioannou:2024tjq. Nuclear experiments only constrain its lower density behavior, i.e., around the nuclear saturation density, $\rho_{\rm{sat}}\equiv 2.7\times 10^{14}\ \rm{g}\ \rm{cm}^{-3}$} (or equivalently, a baryon number density of $n_{\rm{sat}}\equiv$0.16\text{\,}{\mathrm{fm}}^{-3}$\$}). For instance, experiments measuring the neutron-skin thickness of atomic nuclei constrain the symmetry energy at nuclear saturation Birkhan:2016qkr; Thiel:2019tkm; PREX:2021umo; Reed:2021nqk; CREX:2022kgg; Giacalone:2023cet, while heavy-ion collisions uniquely probe isospin-symmetric matter above the nuclear saturation density Danielewicz:2002pu; Li:2008gp; Huth:2021bsp; Sorensen:2023zkk. In addition, nuclear theory calculations provide constraints on isospin-asymmetric matter up to two times the nuclear saturation density Tews:2018kmu; Drischler:2020hwi; Armstrong:2025tza.
Matter at the most extreme densities, up to $6-8n_{\rm{sat}}$, is only realized inside NSs, the remnants of core-collapse supernovae Lattimer:2004sa. Their macroscopic properties, such as mass, radius, quadrupole moment, and tidal deformability, are determined through the Tolman-Oppenheimer-Volkoff (TOV) equations, which require the EOS for closure. Consequently, measurements of NS properties can then be used to infer the EOS Lindblom1992-zb, which is referred to as the inverse problem of NSs. Several such measurements are now available. For instance, precise radio timing of binary pulsar systems determines the masses of NSs Hobbs:2006cd; Demorest:2010bx. Furthermore, pulse-profile modeling of X-ray emissions from hotspots of millisecond pulsars provides joint estimates on the mass and radius of NSs Watts:2019lbs, and has been achieved with the Neutron star Interior Composition ExploreR (NICERgendreau2016neutron; Bogdanov:2019ixe; riley2023xpsi; Riley:2019yda; Miller:2019cac; Riley:2021pdl; Miller:2021qha. Additionally, the tidal deformability of NSs leaves a measurable imprint on gravitational waves emitted during the inspiral and merger of binary neutron star (BNS) systems. This way, GW170817 LIGOScientific:2017vwq; LIGOScientific:2018hze, the first BNS merger observed by Advanced LIGO LIGOScientific:2014pky and Advanced Virgo VIRGO:2014yos, provided the first estimates on NS radii derived from GWs LIGOScientific:2018cki.
Constraining the EOS with input from nuclear physics and astrophysics is commonly done with Bayesian inference Raithel:2017ity; Radice:2017lry; Annala:2017llu; Raithel:2018ncd; Coughlin:2018fis; Capano:2019eae; Miller:2019nzo; Tews:2019cap; Li:2020bre; Legred:2021hdx; Huth:2021bsp; Biswas:2021pvm; Raaijmakers:2021uju; Pang:2021jta; Biswas:2021yge; Essick:2021ezp; Pang:2022rzc; Essick:2020flb; Brandes:2022nxa; Rose:2023uui; Brandes:2024wpq; Rutherford:2024srk; Biswas:2024hja; Huang:2024rfg; Raaijmakers2025; Kashyap:2025cpd; Wu:2025edc (see Ref. Koehn:2024set for a recently compiled overview of EOS constraints). However, current inference methods remain computationally expensive, particularly when directly sampling a set of EOS parameters, as this requires solving the TOV equations on the fly for millions of samples. This computational burden will hinder our ability to extract precise EOS constraints from NS data delivered by future instruments, such as upcoming X-ray telescopes (e.g., eXTP and STROBE-X eXTP:2018anb; Watts:2018iom; STROBE-XScienceWorkingGroup:2019cyd) and the next-generation of ground-based GW detectors (i.e., the Einstein Telescope Punturo:2010zza; ET:2019dnz; ETScienceCase2020; ETDesignUpdate2024 and Cosmic Explorer Evans:2021gyd). For example, next-generation GW observatories are projected to detect approximately $\mathcal{O}(10^{5})$ BNS mergers annually ET:2019dnz; Piro:2021oaa. These detectors will measure NS radii potentially with an unprecedented accuracy of 1010010-100 meters Chatziioannou:2021tdi; Puecher:2023twf; Iacovelli:2023nbv; Huxford:2023qne; Walker:2024loo. This increase in both the quantity and quality of observations highlights the need for more efficient inference methods for the inverse problem of NSs. Ideally, such methods should be capable of solving the TOV equations on the fly to ensure accurate and robust EOS constraints.
One possible solution, thoroughly explored in past works, is to resort to machine-learning surrogates. For instance, previous works have accelerated the TOV solver by replacing it with machine learning-based approximations using regression methods Richter:2023zec; Imam:2023ngm, neural networks Ferreira:2019bny; Thete:2022drz; Tiwari:2024jui; Reed:2024urq; Magnall:2024ffd; DiClemente:2025pbl; Somasundaram:2024ykk, support vector machines Ferreira:2019bny, Gaussian processes and reduced basis methods Reed:2024urq, and dynamic mode decomposition Lalit:2024vmu (see Ref. Zhou:2023pti for a review). Other works, rather than relying on traditional sampling techniques, use machine learning models to directly recover the EOS from NS data, e.g., with neural networks Fujimoto:2017cdo; Fujimoto:2019hxv; Morawski:2020izm; Traversi:2020dho; Krastev:2021reh; Fujimoto:2021zas; Soma:2022qnv; Soma:2022vbb; Farrell:2022lfd; Han:2022sxt; Ferreira:2022nwh; Carvalho:2023ele; Soma:2023fyi; Krastev:2023fnh; Wu:2023npy; Chatterjee:2023ecc; Ventagli:2024xsh, normalizing flows Morawski:2022aud; Brandes:2024vhw; Hu:2024oen, transformers Goncalves:2022smd, principal component analysis Patra:2023jbz, and symbolic regression Imam:2024xqg; Patra:2025sak; Patra:2025xtd; Bejger:2025hbq. However, training these surrogate models can take on the order of hours McGinn:2024nkd to weeks Hu:2024oen (depending on the model complexity and hardware used for training) and requires additional modeling efforts. Moreover, we are often interested in comparing inferences with different parameterizations or prior assumptions. Since emulators must be retrained whenever these change, their computational speed-up comes at the cost of reduced flexibility in inference.
In this work, we propose an alternative solution to efficiently solve the inverse problem of NSs without resorting to machine-learning emulators. In particular, we leverage differential programming and hardware acceleration with graphical processing units to infer the EOS from NS properties in two ways. First, we make Bayesian inference of the EOS from NS data in high-dimensional EOS parameter spaces computationally efficient. Second, we introduce a gradient-based optimization scheme that can recover the parameters of the EOS underpinning a given NS mass-radius curve. While the former method can take observational uncertainties into account, the latter can be used to improve our understanding of how accurate knowledge of NSs propagates into constraints on the EOS.
This paper is organized as follows. In Sec. II, we provide an overview of the EOS parametrization used in this work, and discuss Bayesian inference and our gradient-based optimization scheme. In Sec. III, we validate the results obtained with our Bayesian inference code and investigate its scaling as a function of the dimensionality of the EOS parameter space. Moreover, we demonstrate how our framework can deliver insights into the inverse mapping from NS data to EOS parameters. Finally, we discuss our results in Sec. IV and conclude in Sec. V.
II Methods
II.1 Equation of state
Several approaches exist to parametrize the EOS without resorting to a specific physical description of supra-nuclear matter. These include piecewise polytropes Read:2008iy; Steiner:2010fz; Hebeler:2013nza; Raithel:2016bux; OBoyle:2020qvf, the spectral representation Lindblom:2010bb; Lindblom:2012zi; Lindblom:2018rfr; Fasano:2019zwm; Lindblom:2022mkr, a speed-of-sound parametrization Tews:2018iwm; Greif:2018njt, and non-parametric models based on Gaussian processes Landry:2018prl; Essick:2019ldf; Landry:2020vaw or neural networks Han:2021kjx; Li:2025obt. In this work, we adopt the parametrization of Ref. Koehn:2024set.
At the lowest densities, up to $0.5n_{\rm{sat}}$, the crust EOS is fixed to the model from Ref. Douchin:2001sv. At densities just above the crust-core transition, we assume that matter consists solely of nucleonic degrees of freedom in $\beta$-equilibrium and use a metamodel (MM) parametrization of the EOS Margueron:2017eqc; Margueron:2017lup; Somasundaram:2020chb. This parametrizes the energy per nucleon
[…]
where $n$ represents the baryon number density and $\delta=(n_{n}-n_{p})/n$} is the asymmetry parameter, $n_{n}$ and $n_{p}$ denote the neutron number density and the proton number density, respectively. The isoscalar (saturation) energy $e_{\rm{sat}}$ and the isovector (symmetry) energy $e_{\rm{sym}}$ are defined as a series expansion in $x=(n-n_{\rm{sat}})/3n_{\rm{sat}}$. The Taylor expansion of the former is given by
[…]
Throughout this work, we fix $E_{\rm{sat}}=-$16\text{\,}\mathrm{MeV}$\$ and $n_{\rm{sat}}=0.16\ \rm{fm}^{-3}$. The expansion of the symmetry energy is given by
$\begin{split}e_{\rm{sym}}(n)&=E_{\rm{sym}}+L_{\rm{sym}}x+\frac{1}{2}K_{\rm{sym}}x^{2}\\ &+\frac{1}{3!}Q_{\rm{sym}}x^{3}+\frac{1}{4!}Z_{\rm{sym}}x^{4}+\dots\,.\end{split}$ (3)
The expansion coefficients in Eqs. (2) and (3) are referred to as the nuclear empirical parameters. The MM parametrization allows us to incorporate constraints from nuclear-physics experiments or compare them with similar constraints obtained from astrophysical observations. For instance, the slope of the symmetry energy $L_{\rm{sym}}$ is correlated with both the neutron-skin thickness as well as the radii of NSs Typel:2001lcw; Lattimer:2015nhk; Russotto:2023ari.
However, this parametrization, describing nucleonic degrees of freedom, is inadequate to model the higher-density regime probed by NSs, as new degrees of freedom might become relevant Lattimer:2006qiu; Oertel:2016xsn; Kumar:2023lhv, possibly with a phase transition Baym:2017whm; Pang:2020ilf; Bogdanov:2022faf; Mondal:2023gbf; Essick:2023fso; Prakash:2023afe; Huang:2025vfl. Therefore, we use a phenomenological and model-agnostic approach at higher densities. Specifically, we use a speed-of-sound extension scheme (CSE) parametrization above a breakdown density $n_{\rm{break}}$, which is varied freely in our parametrization within $1-2\ n_{\rm{sat}}$ Tews:2018iwm, unless stated otherwise. At densities above $n_{\rm{break}}$, we parametrize the speed-of-sound $c_{s}$ using a non-uniform grid of points in the $(n,c_{s}^{2})$-plane and construct the full speed-of-sound profile with linear interpolation. We extend this CSE to $25n_{\rm{sat}}$, the end-point of our EOSs, for which $c_{s}^{2}$ is also sampled. The speed-of-sound curve from the CSE parametrization is then matched to that of the MM parametrization with cubic splines. The EOS can then be derived from the speed-of-sound curve, see Ref. Somasundaram:2021clp.
We refer to the parameters of our parametrization, encompassing the NEPs and the CSE grid points, as $\bm{\theta}_{\rm{EOS}}$.
II.2 Neutron star observables
The macroscopic properties of non-spinning, isolated, spherically symmetric NSs are determined through the TOV equations Tolman:1939jz; Oppenheimer:1939ne, the general-relativistic equations determining hydrostatic equilibrium [^1] We refer readers to Ref. glendenning2010special for a didactic derivation of the TOV equations.:
$\frac{\mathop{}\!\mathrm{d}p}{\mathop{}\!\mathrm{d}r}$ $=-\frac{\varepsilon(r)m(r)}{r^{2}}\frac{\left[1+\frac{p(r)}{\varepsilon(r)}\right]\left[1+\frac{4\pi r^{3}p(r)}{m(r)}\right]}{\left[1-\frac{2m(r)}{r}\right]}\,$, (4)
$\frac{\mathop{}\!\mathrm{d}m}{\mathop{}\!\mathrm{d}r}$ $=4\pi r^{2}\varepsilon(r)\,$, (5)
where $\varepsilon$ is the energy density, $r$ is the radial coordinate from the center of the NS and $m(r)$ is the mass contained in a sphere of radius $r$. Here, we employ units where $G=c=1$. The system of equations is closed by the EOS, which links pressure and energy density. Given a central pressure $p_{c}$ of the NS, the TOV equations can be integrated outwards, with boundary conditions m(r=0)=0{m(r=0)=0} and $p(r=0)=p_{c}$}, until we reach the boundary of the star where p(r=R)=0{p(r=R)=0}. This gives the star’s radius $R$ and gravitational mass $M\equiv m(R)$}. Repeating this procedure for various $p_{c}$ provides the mass-radius relation of NSs for that EOS.
When an NS is placed in a static, external quadrupolar tidal field $\mathcal{E}_{ij}$, it develops a quadrupole moment $Q_{ij}$ Hinderer:2009ca; Damour:2009vw; Damour:2012yf. This induced deformation is described by the $\ell=2$ eigenfunctions of the oscillations and, to linear order, takes the form $Q_{ij}=-\lambda\mathcal{E}_{ij}$, where $\lambda$ is related to the second tidal Love number $k_{2}$ through $k_{2}=\frac{3}{2}\lambda R^{-5}$. We refer readers to Refs. Hinderer:2009ca; Damour:2009vw; Damour:2012yf for further details. These tidal deformations occur, for instance, in BNS systems, where the gravitational field of each NS induces the tidal effects deforming its companion. This leaves an imprint on the phase of the GWs emitted during inspiral. In particular, the GW emission depends, at leading order, on each star’s dimensionless tidal deformability parameter $\Lambda$, which is defined as
$\Lambda=\frac{2}{3}k_{2}\frac{R^{5}}{M^{5}}\,.$ (6)
II.3 Differentiable programming
Differentiable programming is a programming paradigm that enables the end-to-end differentiation of computer programs Baydin:2015tfa; sapienza2024differentiableprogrammingdifferentialequations; blondel2024elementsdifferentiableprogramming. The derivative is computed with automatic differentiation, which exploits the fact that all numerical calculations are composed of elementary operations. The derivatives of these operations are accumulated during runtime to efficiently evaluate the gradient of a program.
As a result, the parameters used in a program can easily be optimized using, for instance, gradient descent. If the program expresses a function $\mathcal{L}(\bm{\theta})$, then gradient descent iteratively adapts the parameters $\bm{\theta}$ along the gradient of the objective function:
$\bm{\theta}^{(i+1)}\leftarrow\bm{\theta}^{(i)}\pm\gamma\nabla\mathcal{L}(\bm{\theta}^{(i)})\,$, (7)
where the hyperparameter $\gamma$ is often referred to as the learning rate and the plus sign (minus sign) is used when maximizing (minimizing) the objective function $\mathcal{L}$.
In this work, we use jax frostig2018compiling; kidger2021on, which provides the aforementioned automatic differentiation capabilities in Python. Moreover, jax can compile Python code and execute it on hardware accelerators, such as a GPU. To exploit these features, we have developed a Python package implementing jax-based EOS code and a TOV solver, called jester [^2] https://github.com/nuclear-multimessenger-astronomy/jester. With jester, we can efficiently evaluate the derivative of any numerical function that depends on the EOS parameters, even if it involves solving the TOV equations to obtain the macroscopic properties of the NS as an intermediate step.
While the TOV boundary condition (pressure reaching zero) appears non-differentiable, our implementation circumvents this challenge through two key techniques. First, we parameterize the TOV equations using enthalpy $h$ as the integration variable (similar to, e.g., Ref. Lindblom1992-zb), which transforms the stopping condition from p(r)=0{p(r)=0} to the smooth endpoint h=0{h=0}. Second, we pre-allocate arrays for a maximum number of integration steps and employ bounded while loops to populate the solution arrays, after which jax operations mask and truncate the results to extract only the physically meaningful portion. A similar strategy is employed to truncate the NS solutions at the TOV mass. Since these operations rely entirely on jax primitives, the entire code supports automatic differentiation.
In the following sections, we introduce how we leverage automatic differentiation in gradient-based samplers and a novel gradient-based inversion scheme for NSs.
II.4 Bayesian inference
The inverse problem of NSs (i.e., inferring the EOS from NS data) is often tackled with Bayesian inference, where the goal is to infer the posterior distribution $P(\bm{\theta}_{\rm{EOS}}|d)$ of the EOS parameters given a dataset $d$, representing either NS observations or constraints from nuclear physics. This is achieved by using Bayes’ theorem
$P(\bm{\theta}_{\rm{EOS}}|d)=\frac{P(d|\bm{\theta}_{\rm{EOS}})P(\bm{\theta}_{\rm{EOS}})}{P(d)}\,$, (8)
where $P(d|\bm{\theta}_{\rm{EOS}})$ is the likelihood function, $P(\bm{\theta}_{\rm{EOS}})$ the prior distribution and $P(d)$ the Bayesian evidence.
[…]
It has previously been shown that flowMC generates more effective samples than an MCMC algorithm that only uses a gradient-based sampler Gabrie:2021tlu. In this work, we use the Metropolis-adjusted Langevin algorithm (MALA) for the gradient-based sampler grenander1994representations. We have not tested our setup with the NUTS sampler, a specific implementation of Hamiltonian Monte Carlo (HMC) Betancourt:2017ebh; Hoffman:2011ukg. However, MALA is computationally cheaper to run (see, e.g., Ref. Neural_Langevin_Dynamical_Sampling for runtime comparisons on some numerical experiments) and gives satisfactory results in terms of effective sample size.
II.5 Equation of state constraints
[…]
II.5.1 Chiral effective field theory
[…]
$P(\bm{\theta}_{\rm{EOS}}|d_{\chi\rm{EFT}})\propto\exp\left(\int_{0.75n_{\rm{sat}}}^{n_{\rm{break}}}\frac{\log f(p(\bm{\theta}_{\rm{EOS}};n),n)}{n_{\rm{break}}-0.75n_{\rm{sat}}}\mathop{}\!\mathrm{d}n\right)\,.$ (9)
We utilize the pressure band from Ref. Tews:2018kmu, which extended calculations for pure neutron matter to $\beta$-equilibrium, making it applicable to NS EOS.
The approach used in this work neglects the correlations between pressures. However, we follow the same approach as Ref. Koehn:2024set in order to fully match their setup for our comparison in Sec. III.1.
II.5.2 Mass measurements of pulsars
Mass measurements of the heaviest known pulsars, obtained by evaluating relativistic effects in radio timing, provide a lower bound for the maximal mass of an NS, also referred to as the TOV mass $M_{\rm{TOV}}$. Here, we consider three massive pulsars: PSR J1614-2230 Demorest:2010bx; Shamohammadi:2022ttx with a mass of $1.937\pm 0.014\ M_{\odot}$, PSR J0348+0432 with a mass of $2.01\pm 0.04\ M_{\odot}$ Antoniadis:2013pzd [^3] Note that recent work Saffer:2024tlb argues the mass to be $\sim 1.8M_{\odot}$. However, in this work, we stick to the mass estimate determined previously, to allow for a direct comparison to Ref. Koehn:2024set. and PSR J0740+6620 with a mass of $2.08\pm 0.07\ M_{\odot}$ (but see also Refs. NANOGrav:2019jur; NANOGrav:2023hde). All values are quoted at the $68\%$ credible level. Radio timing measurements $d_{\rm{radio}}$ give a posterior on the pulsar mass $P(M|d_{\rm{radio}})$, which is well approximated by a normal distribution with the mean and standard deviation values quoted above. For each pulsar, the likelihood on the EOS parameters is then
$P(\bm{\theta}_{\rm{EOS}}|d_{\rm{radio}})\propto\frac{1}{M_{\rm{TOV}}}\int_{0}^{M_{\rm{TOV}}}P(M|d_{\rm{radio}})\mathop{}\!\mathrm{d}M\,$, (10)
where the EOS dependence is implicit in $M_{\rm{TOV}}$ and we assumed a uniform distribution on the masses.
II.5.3 NICER
Pulse profile modeling of NICER data has provided mass-radius posteriors for various pulsars Riley:2019yda; Miller:2019cac; Riley:2021pdl; Miller:2021qha; Choudhury:2024xbk; Salmi:2024bss. In this work, we limit ourselves to those of PSR J0030+0451 and PSR J0740+6620. These analyses have later been refined with more accumulated data and advanced models for the hotspots and instrument response Vinciguerra:2023qxq; Salmi:2024aum; Dittmann:2024mbo. We use posterior samples on the masses and radii from the public data releases NICER_posterior_samples_Zenodo_J0030_Riley; NICER_posterior_samples_Zenodo_J0030_Miller; NICER_posterior_samples_Zenodo_J0740_Salmi; NICER_posterior_samples_Zenodo_J0740_Dittmann. Their distributions $P(M,R|d_{\rm{NICER}})$ are approximated by a Gaussian kernel density estimate. In our inference, we sample the mass $M$ of each pulsar from a uniform prior distribution and compute the likelihood as
$P(\bm{\theta}|d_{\rm{NICER}})\propto P(M,R(\bm{\theta}_{\rm{EOS}};M))|d_{\rm{NICER}})\,$, (11)
where the parameters $\bm{\theta}$ consist of the EOS parameters and the pulsar masses.
II.5.4 GW170817
Finally, we consider constraints from tidal deformability measurements of GW170817.444The second observed BNS candidate so far, GW190425 LIGOScientific:2020aai, does not significantly constrain the EOS due to its lower signal-to-noise ratio and larger component masses, and therefore will not be considered in this work. For this, one first has to determine the posterior distributions on the source parameters $\bm{\theta}_{\rm{GW}}$ of GW170817 Veitch:2009hd; Veitch:2014wba; Romero-Shaw:2020owr. Here, we take the posterior samples produced by the analysis of Ref. Wouters:2024oxj, which obtained the posterior using Jim Edwards:2023sak; Wong:2023lgb; Wouters:2024oxj, a GW inference toolkit implemented in jax able to perform fast Bayesian inference on GW signals. The samples are publicly available at Ref. Jim_GW170817_posterior_samples.
From the posterior on the detector-frame chirp mass, the mass ratio, and the luminosity distance, we infer the source-frame component masses $M_{i}$ using the linear Hubble law, where $i=1,2$ labels the two components of the binary. For the inference on the EOS, we then consider the marginalized posterior distribution on $M_{i}$ and the tidal deformabilities $\Lambda_{i}$. Contrary to the case of the NICER mass-radius posteriors, a Gaussian kernel density estimate provides a poor approximation of this distribution. Therefore, we train [^5] Training this NF takes less than $\lesssim 15$ minutes on an NVIDIA A100 GPU using FlowJAX ward2023flowjax. an NF to obtain a tractable and more accurate approximation of the marginal posterior density, which we denote by $P_{\rm{NF}}(M_{i},\Lambda_{i}|d_{\rm{GW}})$. During inference, we sample the binary component masses $M_{i}$ and use the likelihood function
$P(\bm{\theta}|d_{\rm{GW}})\propto P_{\rm{NF}}(M_{i},\Lambda_{i}(\bm{\theta}_{\rm{EOS}};M_{i})|d_{\rm{GW}})\,$, (12)
where $\bm{\theta}$ consists of the EOS parameters and source-frame component masses of GW170817.
II.6 Gradient-based inversion
To assess how detailed knowledge on NS properties propagates to the EOS and the parametrization we employ for it, we have to invert a complete mass-radius relation $R(M)$ or mass-tidal deformability relation $\Lambda(M)$. Here, we show that such an inversion can be achieved by employing a gradient-based optimization approach for the EOS parameters, as an alternative to Bayesian inference, as the latter can be quite computationally expensive.
[…]
$\mathcal{L}(\bm{\theta}_{\rm{EOS}})=\frac{1}{N}\sum_{i=1}^{N}\left|\frac{\Lambda_{i}(\bm{\theta}_{\rm{EOS}})-\hat{\Lambda}_{i}}{\hat{\Lambda}_{i}}\right|\,.$ (13)
We note that the loss function is based on mock data coming from the “true” EOS rather than existing NS observations, and therefore is introduced in this work as a proof of concept rather than a full replacement of Bayesian inference.
By starting from a randomized initial position in the EOS parameter space and iteratively performing gradient descent on this loss function, as shown in Eq. (7), we aim to recover a set of EOS parameters that matches the given mass-tidal-deformability relation. The derivative of the loss function is computed with automatic differentiation, as explained in Sec. II.3. In particular, we use the Adam optimizer, which stabilizes the descent with adaptive estimates of the first and second moments of the gradients Kingma:2014vow. Figure 1 illustrates our optimization algorithm. The NS curve of the EOS at the final iteration deviates from the target NS family with an error less than $100$ meters in radius and less than $10$ in tidal deformability for the mass range considered during optimization. We note that we show the optimization run after truncating around $600$ iteration steps for visualization purposes. We expect further improvement in the convergence, especially at lower masses, if the optimizer runs for more iterations. Since the “true” EOS considered in this example comes from the same EOS parametrization as used in the recovery, a perfect recovery of the EOS is theoretically possible. However, in practice, the optimization can get stuck in local minima. This might be more problematic in case the EOS does not follow the same parametric form as used in the recovery. Moreover, incorporating real data with uncertainties will change the loss function of Eq. (13) and the robustness of the recovery procedure. These additional complications will be addressed in future work.
Figure 1: Illustration of the gradient-based inversion scheme recovering a given mass-radius relation of NSs. The color gradient shows the different iterations while performing gradient descent on the objective function of Eq. (13). The gradient is computed with respect to the EOS parameters throughout the TOV equations using automatic differentiation. Top panel: Trajectory of the EOS parameters, showing only $L_{\rm{sym}}$ and $K_{\rm{sat}}$ for simplicity. Middle panel: Mass-radius and mass-tidal deformability curves during the gradient descent, with the target shown in black. Lower panel: Pressure as a function of density curves during optimization.
While preparing our manuscript, Refs. Lindblom:2025kjz; Li:2025obt also introduced a recovery algorithm based on optimization. However, their methods differ from ours in several ways. Ref. Lindblom:2025kjz uses a spectral expansion Lindblom:2024yio and relatively low-dimensional EOS parameter spaces. Our scheme, instead, scales well to high-dimensional parameter spaces by employing gradient descent. Indeed, the illustration shown in Fig. 1 performed the inversion on the complete EOS parametrization described in Sec. II.1, using eight grid points for the CSE, and was able to recover the EOS. Since this parametrization involves $26$ parameters in total, this example demonstrates the applicability of our method in high-dimensional EOS spaces. Ref. Li:2025obt uses a gradient-based inversion scheme, but uses a neural network as EOS parametrization and relies on linear response analysis to compute the gradients. In contrast, we use automatic differentiation, which enables us to compute these gradients for any EOS parametrization.
III Results
III.1 Validation and computational scaling
We first demonstrate that our Bayesian inference methods reproduce previous results, and then investigate the scaling of the methodology as a function of the number of dimensions employed in the parametrization.
In order to validate our method, we compare our results to those of Ref. Koehn:2024set which used an identical EOS parametrization. [^7] The quoted results for Ref. Koehn:2024set are obtained from https://multi-messenger.physik.uni-potsdam.de/eos_constraints/. Moreover, we use the same prior distribution, given in Table 1. We take eight grid points for the CSE parametrization.
Table 1: Table of EOS parameters and their uniform prior distribution ranges, separated in the metamodel and speed-of-sound extension (CSE) parts, with the index $i$ labeling the grid points. Throughout this work, we fix $n_{\rm{sat}}=0.16\ \rm{fm}^{-3}$ and $E_{\rm{sat}}=-16$ MeV.
Metamodel Parameter Prior
$K_{\rm{sat}}$ [MeV] $U(150,300)$
$Q_{\rm{sat}}$ [MeV] U(500,1100)U(-500,1100)
$Z_{\rm{sat}}$ [MeV] U(2500,1500)U(-2500,1500)
$E_{\rm{sym}}$ [MeV] $U(28,45)$
$L_{\rm{sym}}$ [MeV] $U(10,200)$
$K_{\rm{sym}}$ [MeV] U(300,100)U(-300,100)
$Q_{\rm{sym}}$ [MeV] U(800,800)U(-800,800)
$Z_{\rm{sym}}$ [MeV] U(2500,1500)U(-2500,1500)
CSE $n_{\rm{break}}$ $n_{\rm{sat}}]$ $U(1,2)$
$n^{(i)}$ $n_{\rm{sat}}]$ $U(n_{\rm{break}},25)$
$(c_{s}^{(i)}/c)^{2}$ $U(0,1)$
In Table 2, we compare the $95\%$ credible intervals of the inferred radius of a $1.4M_{\odot}$ NS, $R_{1.4}$. The first rows of the table correspond to inferences employing the constraints discussed in Sec. II.5 individually. The bottom row shows the result when considering all constraints jointly in a single Bayesian inference run. The table demonstrates that our results qualitatively agree with Ref. Koehn:2024set, but show some quantitative differences in the inferred radii. Nevertheless, for all constraints considered here, we do observe quantitative agreement between both approaches for $M_{\rm{TOV}}$, the pressure of the EOS at $3n_{\rm{sat}}$, and the maximal central density inside a NS, $n_{\rm{TOV}}$, as shown by Table 3 in Appendix A.
[…]
Table 2: Comparison of inferred $R_{1.4}$ ($95\%$ credible interval) in kilometers between Ref. Koehn:2024set and this work for different EOS constraints. The bottom row refers to the posterior using all listed constraints jointly.
$R_{1.4}$ [km]
Constraint Ref. Koehn:2024set This work
$\chi_{\rm{EFT}}$ $12.11_{-3.39}^{+1.69}$ $12.59_{-3.51}^{+2.24}$
Radio timing $13.70_{-2.17}^{+1.41}$ $13.71_{-1.88}^{+1.19}$
PSR J0030+0451 $13.17_{-2.24}^{+1.65}$ $13.48_{-2.15}^{+1.42}$
PSR J0740+6620 $13.39_{-1.72}^{+1.57}$ $13.79_{-1.73}^{+1.26}$
GW170817 $11.90_{-1.74}^{+1.78}$ $12.24_{-2.18}^{+2.03}$
All $12.26_{-0.91}^{+0.80}$ $12.62_{-1.11}^{+0.97}$
After validating that our framework agrees with previously published results, we now assess the runtime and its scaling as a function of the dimensionality of the EOS parameter space. Using a single NVIDIA H100 GPU, the inferences listed in Table 2 were performed under $1$ hour of wall time. Solving the TOV equations to obtain the macroscopic properties of NSs ($100$ masses and their corresponding radii, and tidal deformabilities) for a single proposal takes around $\sim 0.3$ milliseconds, which is on par with previous works relying on machine learning surrogates (see, e.g., Refs Reed:2024urq; Magnall:2024ffd; Tiwari:2024jui for recent examples using neural networks), however, without the need of pretraining a machine learning model in our case.
In the future, more precise observations of NSs will become available, which could require a more flexible EOS parameterization to avoid biases. This motivates us to extend the parameter space by adding more CSE grid points and assess the impact on the scaling of our sampling efficiency. Specifically, we perform Bayesian inference jointly on all EOS constraints considered in this work by testing CSE grids with $10$, 2020, and $30$ points, respectively. Figure 2 shows the runtime per effective sample size (ESS), estimated with Geyer’s initial monotone sequence criterion geyer_SSE_ref1; geyer_SSE_ref2. The axis on the right converts this number to an estimate of the total wall time required in case one requires $5000$ effective samples. We perform the scaling runs on two different GPU architectures, namely an NVIDIA A100 and NVIDIA H100, and show the average and standard deviation over all the EOS parameters. The parametrization with the highest number of grid points increases the runtime roughly by a factor $4$ to produce a similar number of effective samples. Therefore, our inferences still only require a couple of hours to provide accurate posteriors on the EOS, even when using a more flexible parametrization.
Figure 2: Scaling of the runtime divided by effective sample size (ESS) as a function of the number of EOS parameters when varying the number of grid points for the CSE parameterization, for two different GPU architectures. The right axis converts this number to total wall time to obtain $5000$ effective samples.
III.2 Constraining $n_{\rm{break}}$ with neutron star data
The EOS parametrization employed in this work takes the breakdown density $n_{\rm{break}}$ of the metamodel as a parameter that can be varied freely, after which the CSE part of the parametrization is used for the high-density EOS. This density is an indicator of the densities for which a simple nucleonic description of the EOS is not sufficient anymore. We can use our framework to constrain $n_{\rm{break}}$ using NS data and directly probe this breakdown with astrophysical information. To this end, we perform Bayesian inference as explained above with two key modifications. First, we use a wider, agnostic prior on $n_{\rm{break}}$, namely a uniform distribution between $1n_{\rm{sat}},4n_{\rm{sat}}]$. The priors on the other MM and CSE parameters are identical to those shown in Tab. 1. Second, we only consider EOS constraints involving NS data, i.e., we do not use the constraint from $\chi_{\rm{EFT}}$. Figure 3 shows the marginal posterior distribution of $n_{\rm{break}}$. We infer a breakdown density of $n_{\rm{break}}=2.47_{-1.02}^{+1.53}\ n_{\rm{sat}}$ at the $90\%$ credible level. Hence, NS data prefers a smooth nucleonic EOS up to 2-3 times the nuclear saturation density, a range within the reach of microscopic nuclear theory Essick:2020flb. While $n_{\rm{break}}$ is not tightly constrained by the data considered in this work, future detectors providing more NS data can potentially improve its precision. Nevertheless, our result serves as proof of concept that the metamodel breakdown density can be determined on-the-fly with Bayesian inference.
We point out that Ref. Biswas:2020puz similarly inferred this transition density. However, comparing the two results is complicated, as they use different parameterizations and prior choices. In particular, we sample the $n_{\rm{break}}$ parameter on the fly during inference, whereas Ref. Biswas:2020puz infers this parameter by considering Bayes factors of different runs with fixed $n_{\rm{break}}$. Moreover, we extend the MM parametrization to higher orders and consider a speed-of-sound parametrization at high densities, rather than piecewise polytropes.
Figure 3: Posterior on $n_{\rm{break}}$ of the metamodel parametrization inferred from a uniform prior between $1n_{\rm{sat}}$ to $4n_{\rm{sat}}$ (other priors are given in Tab. 1) and given the NS observations from Tab. 2.
III.3 Constraints on nuclear empirical parameters
Currently available NS observations do not yet tightly constrain the EOS, as demonstrated by the previous section. In this section, we instead consider a hypothetical scenario where observations of NSs are pushed to extreme limits and uncertainties on the measured radii and tidal deformabilities can be ignored. We aim to obtain a qualitative estimate of how highly accurate knowledge of the properties of NSs translates into constraints on the NEPs, since such constraints can then be compared to those obtained from nuclear experiments. For simplicity, we restrict our parametrization to the metamodel part and no longer use the CSE at high densities. We enforce all these metamodel EOSs to be thermodynamically stable as well as to respect causality. Otherwise, the EOSs are truncated at the maximal central density allowed in an NS.
We generate a mock “true” EOS [^8] Here, we did not consider agreement between this mock EOS and the constraints from perturbative QCD at high densities, as this merely serves as a proof of concept of the methodology. with a TOV mass of roughly $2.0M_{\odot}$ and aim to recover its EOS parameters from its $M$-$\Lambda$ curve using the gradient-based optimization scheme discussed in Sec. II.6. In particular, the inversion is performed with perfect knowledge of the $M$-$\Lambda$ curve, as discussed in detail in Sec. II.6. We leave a full study which considers Bayesian inference using the constraints from Tab. 2for future work. We perform multiple such recoveries where we change the number of NEPs that are varied freely in the recovery, starting with only the first order parameter $L_{\rm{sym}}$. In subsequent recovery runs, we also allow the optimizer to explore the higher-order parameters, each time adding the parameters of the next order of the Taylor expansions of Eqs. (2)-(3). For simplicity, we do not vary the $E_{\rm{sym}}$ parameter, since it has a small impact on the EOS at high densities, and we wish to focus on the parameters entering at higher orders of the Taylor expansion. For each recovery, we start the optimizer from $100$ random EOS samples, sampled from the priors shown in Table 1, and perform at most $1000$ iterations with a fixed learning rate of 11.999The runs are performed on an AMD EPYC 7543 VM CPU and take at most a few minutes per recovery. For the final result, we keep only those EOS samples that within the $1\ M_{\odot},2\ M_{\odot}]$ mass range deviate less than $100$ meters in radius from the “true” $M$-$R$ curve and less than $10$ in tidal deformability from the “true” $M$-$\Lambda$ curve.
Figure 4 shows the slope of the symmetry energy $L_{\rm{sym}}$ for these recovered EOSs. The range of the recovered $L_{\rm{sym}}$ values significantly depends on the number of NEPs that are allowed to vary. When $L_{\rm{sym}}$ is the only degree of freedom, we recover the value of the “true” EOS, with a small width of around $0.5\ \rm{MeV}$, as expected. However, when more NEPs are varied, this range increases. Nevertheless, the middle and right panels of Fig. 4 demonstrate (for the case where all NEPs are free parameters) that the corresponding EOSs and NS configurations agree well with the true EOS. Therefore, from our gradient-based optimization approach, we conclude that, while accurate observations of NSs allow us to tightly constrain the EOS itself, the corresponding constraints on the NEPs suffer from a degeneracy in our metamodel parametrization, which was already hinted at by Refs. Krastev:2018nwr; Xie:2019sqb; Xie:2020tdo; Mondal:2021vzt; Iacovelli:2023nbv. Therefore, fitting relations between $L_{\rm{sym}}$ and $R_{1.4}$ previously proposed in the literature, e.g., Ref. Raithel:2020vvg, will need to be assessed carefully in light of the assumptions regarding the parametrization used to obtain them. Finally, we note that this degeneracy is expected to be more pronounced if we include the CSE parametrization at higher densities, as this would further diminish the impact of the individual NEPs at the higher-density regime of the EOS to which NSs are most sensitive.
Figure 4: Recovery of an EOS in the metamodel parametrization from the $M$-$\Lambda$ curve using the gradient-based optimization algorithm as a function of the number of NEPs being varied. The left panel shows the recovered range of $L_{\rm{sym}}$ when increasingly more NEPs at higher orders in the Taylor expansion are varied freely. For the recovery in which all NEPs were varied freely, we show the recovered EOS samples in the middle and right panels. The middle panel shows the pressure as a function of density, while the right panel shows the mass-radius profile of NSs, color-coded by their $L_{\rm{sym}}$ value. The grey bands show mass ranges not considered by the loss function in Eq. (13). The EOSs shown in this figure deviate from the target by less than $100$ meters in radius and less than $10$ in dimensionless tidal deformability across the mass range considered in the loss function of Eq. (13).
IV Discussion
IV.1 Equation of state degeneracies
Our results from Sec. III.3 show that the inverse problem of NSs is sensitive to degeneracy in our metamodel parametrization. While this was shown using our gradient-based optimization scheme, future work will assess if these degeneracies remain when using traditional stochastic sampling algorithms. A similar degeneracy was already observed in previous works that used stochastic samplers, e.g., Refs. Krastev:2018nwr; Xie:2019sqb; Xie:2020tdo; Mondal:2021vzt; Iacovelli:2023nbv. However, our work extends these studies by simultaneously considering (i) higher-order NEPs, (ii) a wide range of NS masses, (iii) high accuracy in the recovered EOSs, highlighting the problem’s relevance for future detectors, and (iv) the dependence as a function of the parametrization. This degeneracy is expected since the MM expansion from Eq. (2) - (3) is performed at $n=n_{\rm{sat}}$, whereas NSs observations probe the EOS at densities of a few $n_{\rm{sat}}$, hinting that other parameterizations which avoid a Taylor expansion at a specific density could potentially alleviate this issue. While our analysis in Sec. III.3 served as a theoretical exploration of this degeneracy, we expect that a full Bayesian analysis employing future constraints from nuclear theory and experiments on the NEPs, especially for the higher-order parameters of the MM, alleviates this issue.
We expect the level of degeneracy to be different for different EOS models, and we will explore such model dependencies in future work. Furthermore, we will also investigate the potential introduction of similar systematic effects originating from certain aspects neglected in this study, such as the crust treatment (see, e.g., Refs. Fortin:2016hny; Gamba:2019kwu; Ferreira:2020wsf; Rather:2020gja; Suleiman:2021hre; Canullan-Pascual:2025smm), and consider unified EOS models Davis:2024nda as alternative.
IV.2 Next-generation detectors
Future detectors will generate an unprecedented volume of NS observations. Each event provides valuable constraints on the EOS, yet, when analyzed with existing methods, at substantial economic and environmental costs Hu:2024mvn. Our work addresses this pressing data analysis challenge by employing differentiable programming and GPU accelerators. Using jax, analyzing a GW signal from a BNS merger to obtain masses and tidal deformabilities can be completed within minutes Wouters:2024oxj, after which we obtain posteriors on the EOS in less than $1$ hour on an NVIDIA H100 GPU. Crucially, we do not require pre-trained machine-learning emulators of the TOV equations, thereby ensuring accuracy in the results and further reducing computational costs. Therefore, our methods are a key advancement towards handling the data analysis problems offered by future detectors.
[…]
IV.3 Extension to other analyses
In this work, we only focused on inferring the EOS from NS observations. However, other inferences require jointly modeling the EOS together with additional effects or parameters. Naturally, these analyses can therefore benefit from the acceleration offered by the methods presented in this work.
For instance, we can infer the EOS simultaneously with cosmological parameters or population models (see, e.g., Refs. Ghosh:2024cwc; Magnall:2024ffd; Golomb:2024lds). Moreover, our method can be extended to jointly infer the parameters of modified gravity, either for a specific theory (see Ref. Olmo:2019flu for an overview) or for a parametrized post-TOV formalism Glampedakis:2015sua; Glampedakis:2016pes; daSilva:2024qex. Our gradient-based scheme can be used to explore this extended parameter space and uncover degeneracies between the EOS and deviations from general relativity in observations of NSs Brown:2024joy. Similar degeneracies can arise from other effects, such as phase transitions Raithel:2022aee; Raithel:2022efm; Counsell:2025hcv or dark matter Koehn:2024gal, which can also be constrained with either Bayesian inference or the gradient-based method.
Going beyond extreme matter studies, other analyses in related fields that have to solve ordinary differential equations on the fly during inference can benefit from the methods discussed in this work, such as the evaluation of effective one-body gravitational waveforms Damour:2009zoi, solving the lens equation in strong lensing analyses (see, e.g., Refs. Grespan:2023cpa; Poon:2024zxn), and nuclear reaction networks Reichert:2023xqy.
V Conclusion
The inverse problem of NSs, i.e., inferring the nuclear EOS from NS data, is a computationally challenging task, making it difficult to understand the interplay between nuclear physics and astrophysics. Moreover, the prospect of a large volume of NS observations delivered by future detectors calls for an efficient and scalable way of addressing this inverse problem. In this work, we have presented a solution based on differentiable programming, leveraging the automatic differentiation capabilities and support for hardware accelerators, such as GPUs, from jax. By using gradient-based samplers, we can perform MCMC sampling on high-dimensional EOS parameter spaces in under an hour of wall time on a single NVIDIA H100 GPU without resorting to machine learning emulators for the TOV equations. Moreover, we provide a proof of concept on how the breakdown of the metamodel description can be inferred from NS data on the fly during sampling.
Additionally, our framework introduces a novel optimization scheme that recovers the EOS of a given mass-radius or mass-tidal deformability relation by performing gradient descent on a loss function formulated in terms of the EOS parameters, differentiating through the TOV equations. Using this tool, we have discussed a degeneracy present in our metamodel parametrization, which might affect inferences of nuclear empirical parameters from detailed knowledge of NSs using certain EOS parametrizations. This optimization scheme offers a new way to investigate the relationship between the EOS and the properties of NSs, and future work will investigate its robustness when the parametrization of the EOS is misspecified or noisy observations are used.
In short, our methods leveraging differentiable programming address the data analysis challenges of future detectors and offer a pragmatic way to investigate the relation between EOS parameterizations and properties of NSs.
Appendix A Method validations
In Table 3, we extend the results of Table 2 and show $M_{\rm{TOV}}$, the pressure at $3n_{\rm{sat}}$, $p(3n_{\rm{sat}})$, and the maximal central density $n_{\rm{TOV}}$ at the core of an NS, between Ref. Koehn:2024set and this work for the EOS constraints discussed in Sec. II.5.
Table 3: Comparison of inferred $M_{\rm{TOV}}$, the pressure at $3n_{\rm{sat}}$, $p(3n_{\rm{sat}})$, and the maximal central density $n_{\rm{TOV}}$ at the core of an NS, between Ref. Koehn:2024set and this work for various EOS constraints. Uncertainties are quoted at the $95\%$ credible level.
Constraint $_{\rm{TOV}} [MM_{\odot}$] $p(3n_{\rm{sat}})$ [MeV fm-3] $n_{\rm{TOV}}$ $n_{\rm{sat}}$]
Ref. Koehn:2024set This work Ref. Koehn:2024set This work Ref. Koehn:2024set This work
$\chi_{\rm{EFT}}$ $2.05_{-1.16}^{+1.08}$ $2.03_{-0.97}^{+1.03}$ $\phantom{0}69_{-53}^{+186}$ $\phantom{0}72_{-65}^{+165}$ $6.51_{-3.11}^{+10.7}$ $6.53_{-4.09}^{+9.26}$
Radio timing $35_{-0.29}^{+0.73} 2.2$$20_{-0.26}^{+0.39}$ $111_{-49}^{+140}$ $\phantom{0}97_{-49\phantom{0}}^{+65\phantom{0}}$ $5.51_{-1.66}^{+1.89}$ $5.68_{-1.74}^{+1.63}$
PSR J0030+0451 $2.16_{-0.71}^{+0.83}$ $2.19_{-0.76}^{+0.78}$ $\phantom{0}89_{-46}^{+143}$ $\phantom{0}94_{-62}^{+135}$ $5.62_{-1.91}^{+4.44}$ $5.60_{-2.59}^{+3.65}$
PSR J0740+6620 $2.34_{-0.32}^{+0.65}$ $2.38_{-0.42}^{+0.70}$ $07_{-40}^{+125} 11$$18_{-59}^{+150}$ $5.34_{-1.61}^{+1.63}$ $5.16_{-2.04}^{+1.72}$
GW170817 $1.90_{-0.41}^{+0.71}$ $1.93_{-0.48}^{+0.73}$ $\phantom{0}59_{-25\phantom{0}}^{+91\phantom{0}}$ $\phantom{0}62_{-42\phantom{0}}^{+109}$ $7.37_{-2.72}^{+3.64}$ $7.24_{-3.21}^{+3.77}$
All $25_{-0.22}^{+0.42} 2.2$$23_{-0.23}^{+0.41}$ $\phantom{0}90_{-31\phantom{0}}^{+71\phantom{0}}$ $\phantom{0}92_{-42\phantom{0}}^{+70\phantom{0}}$ $5.92_{-1.38}^{+1.35}$ $5.92_{-1.49}^{+1.38}$
Figure 5: Comparison of prior distributions between Ref. Koehn:2024set (Koehn+2024) and this work on the EOS quantities shown in Tab. 2 and Tab. 3. Since both use the EOS parametrization described in Sec. II.1, with priors given by Tab. 1, remaining differences are due to our jax implementation.
Data availability
Our EOS code and TOV solver are available open source at nuclear-multimessenger-astronomy/jester. The code used in this study is available at Ref. Github_repo_paper.
Acknowledgments
We thank Frank Snijder, Fabian Gittins, Harsh Narola, Justin Janquart, Anna Watts, Mariska Hoogkamer, Michael Coughlin, Micaela Oertel, and the LIGO-Virgo-Kagra extreme matter community for fruitful discussions and feedback that led to the improvement of this work. T.W. and C.V.D.B. are supported by the research program of the Netherlands Organization for Scientific Research (NWO) under grant number OCENW.XL21.XL21.038. P.T.H.P. is supported by the research program of the Netherlands Organization for Scientific Research (NWO) under grant number VI.Veni.232.021. T.W., P.T.H.P., H.K., H.R., T.D. acknowledge support from the Daimler and Benz Foundation for the project “NUMANJI”. T.D., H.K., H.R. acknowledges support from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. R.S. acknowledges support from the Nuclear Physics from Multi-Messenger Mergers (NP3M) Focused Research Hub which is funded by the National Science Foundation under Grant Number 21-16686. R.S. and I.T. were supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). R.S. and I.T. were also supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) NUCLEI program, and by the Laboratory Directed Research and Development Program of Los Alamos National Laboratory under project number 20230315ER. We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius under project numbers EINF-6587 and EINF-8596. The authors acknowledge the computational resources provided by the LIGO Laboratory’s CIT cluster, which is supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.
AD
automatic differentiation
JIT
just-in-time
PE
parameter estimation
MCMC
Markov chain Monte Carlo
GW
gravitational wave
EM
electromagnetic
CBC
compact binary coalescences
NS
neutron star
KDE
kernel density estimate
NF
normalizing flow
BBH
binary black hole
BNS
binary neutron star
NSBH
neutron star-black hole
EOS
equation of state
EFT
effective field theory
[…]
chiral effective field theory
NEP
nuclear empirical parameter
HIC
heavy-ion collision
$M$
metamodel
CSE
speed-of-sound extension scheme
TOV
Tolman-Oppenheimer-Volkoff
JS
Jensen-Shannon
CPU
central processing unit
GPU
graphical processing unit
TPU
tensor processing unit
ML
machine learning
SNR
signal-to-noise ratio
PSD
power spectral density
NICER
Neutron star Interior Composition ExploreR